The force on a particle is directed along an x axis and given by F = F0(x/x0 - 1) where x is in meters and F is in Newtons. If F0 = 1.6 N and x0 = 2.0 m, find the work done by the force in moving the particle from x = 0 to x = 2x0 m.
Answer #1

F = Fo(X/Xo - 1)

F = 1.6*(x/2 -1)

Work done W = (integral from x = 0 to x = 2xo)(F*dx)

W = (integral from x = 0 to x = 2xo)(1.6*(x/2 -1)*dx)

W =(from x = 0 to x = 2xo)1.6*(x^2/4 - x)(from x = 0 to x = 2xo)

W = 1.6*(4*xo^2/4 - 2xo)-(0-0)

W = 1.6*(xo^2 - 2xo)

W = 1.6*(2^2-2*2) = 0 J